# Analytical bounds for decoy-state quantum key distribution with discrete phase randomization

### v1

@@ IV-B Analytical bounds for MDI QKD with DPR
In the DPR case, a lower bound on the key rate of the MDI-QKD protocol is given by Eq. 17. The variables we need to bound are $Y_{\gamma,kk}^{\mu\mu}=Q_{\gamma,kk}^{\mu\mu}/(p_{k}^{\mu}p_{k}^{\mu})$ and $e_{y,kk}^{b,\mu\mu}$, for $k\in\{0,1\}$, from which $e_{x,kk}^{p,\mu\mu}$ can be bound.
@@
$W_{y,00}^{\mu\mu}$ $\leq W_{y,00}^{\mu\mu U}\equiv\max(Q_{y}^{00}E_{y}^{00}+2\epsilon_{\mu 0},0.5),$ (26)
where $Q_{y}^{00}$ and $Q_{y}^{00}E_{y}^{00}$ are observables, and $W_{\gamma,k}^{\alpha\beta}=e_{\gamma,k}^{b,\alpha\beta}Y_{\gamma,k}^{\alpha\beta}$. An upper bound for the bit error rate is $e^{b,\mu\mu}_{y,00}\leq e^{b,\mu\mu U}_{y,00}=W_{y,00}^{\mu\mu U}/Y_{y,00}^{\mu\mu L}$.
In terms of bounds on $Y_{\gamma,11}^{\mu\mu}$ and $W_{y,11}^{\mu\mu}$, we use Eq. 19 and some mathematical inequalities in Appendix B to obtain
@@
can be calculated from observed values; $G$ is given by $\text{max}(A,B,C)$, where
@@
$\epsilon_{s}=$ $e^{\alpha+\beta}(\sum_{a=2}^{N-1}p_{a}^{\alpha}p_{1}^{\beta}(\epsilon_{\alpha\mu}+\epsilon_{\beta\mu})+\sum_{b=2}^{N-1}p_{1}^{\alpha}p_{b}^{\beta}(\epsilon_{\alpha\mu}+\epsilon_{\beta\mu})$
$+\sum_{a=2}^{N-1}\sum_{b=2}^{N-1}p_{a}^{\alpha}p_{b}^{\beta}(\epsilon_{\alpha\mu}+\epsilon_{\beta\mu})),$
@@
Next, the upper bound on $W_{y,11}^{\mu\mu}$ is calculated from
$\begin{split}W_{y,11}^{\mu\mu}\leq W_{y,11}^{\mu\mu U}=\frac{e^{\alpha+\beta}Q_{y}^{\alpha\beta}E_{y}^{\alpha\beta}-T_{2}-\epsilon_{\bar{s}}}{e^{\alpha+\beta}p_{1}^{\alpha}p_{1}^{\beta}},\end{split}$ (31)
@@
$\begin{split}T_{2}=e^{\alpha+\beta}(p_{0}^{\alpha}Q_{y}^{0\beta}E_{y}^{0\beta}+p_{0}^{\beta}Q_{y}^{\alpha 0}E_{y}^{\alpha 0}-p_{0}^{\alpha}p_{0}^{\beta}Q_{y}^{00}E_{y}^{00}).\end{split}$ (32)
@@ V Simulation Results and Discussion
In our simulations, we consider the scenario in which no eavesdropper (Eve) is present, but system parameters and imperfections determine the observed values in an experiment. In our simulations of the BB84 protocol, the total gain and the corresponding QBER, for an intensity $\alpha$, are, respectively, assumed to be [ma2005practical, panayi2014memory, cao2015discrete]
$Q_{\gamma}^{\alpha}$ $=Y_{0}+1-e^{-\eta\alpha},$
@@
where $Y_{0}=2p_{d}$ denotes the vacuum yield, $p_{d}$ is the dark count rate per pulse, $\eta$ represents the overall channel transmittance, $e_{0}=0.5$ is the error rate associated with vacuum states, and $e_{d}$ denotes the phase misalignment error rate. In our simulations of the MDI QKD protocol, the total gain and the corresponding QBER when Alice sends states with intensity $\alpha$ and Bob sends states with intensity $\beta$ are, respectively, taken as [lo2012measurement, cao2020discrete]
$Q_{\gamma}^{\alpha\beta}=$ $(Y_{00}+1-e^{-\eta\alpha})(Y_{00}+1-e^{-\eta\beta)}),$
$Q_{\gamma}^{\alpha\beta}E_{\gamma}^{\alpha\beta}=$ $Y_{00}(Y_{00}+2-e^{-\eta\alpha}-e^{-\eta\beta})/2$
$+e_{d}(1-e^{-\eta\alpha})(1-e^{-\eta\beta}),$ (34)
where $Y_{00}$ is assumed to be $2p_{d}$. Using the above values as our observables, we can use the techniques introduced in Section III and Section IV to bound the key rate.
Figure 3: Secret key generation rates for the DPR BB84 protocol using numerical (solid lines) and analytical (dashed) methods. The mean number of photons for the signal and decoy states are optimized to get the maximum key rate. The best performing curve is based on the vacuum+weak decoy states and CPR. The remaining curves are generated when considering $D$ phase slices and vacuum+weak decoy states. The simulation parameters are shown in Table I [gobby2004quantum, cao2015discrete].
Figure 4: Bit error rate $e_{y,1}^{b,\mu}$ for the DPR BB84 protocol using numerical (solid) and analytical (dashed) methods. Results for the CPR case are shown as well. For clarity, the curves corresponding to $D<9$ are smoothed to better display the relationship between the bit error rate and either the phase slice number or the communication distance.
Figure 5: The optimized value of $\mu$ for different number of phase slices ranging from $D=5$ to $D=10$, versus distance in the DPR BB84 protocol. A CPR related curve is provided for comparison. Only the numerical results are given here because we get almost the same value of optimized $\mu$ from both methods. All curves are smoothed to mask numerical fluctuations.
@@
Figure 3 presents the secret key rates for the vacuum+weak decoy BB84 protocol, using the numerical method (solid lines) described in Section III-A, by solving the optimization problem in Section II-C, Approximate Formulation, and the analytical method (dashed lines) outlined in Section IV-A. The corresponding bit error rates are shown in Fig. 4. The simulation parameters are shown in Table I [gobby2004quantum, cao2015discrete]. During the optimization process for both numerical and analytical methods, the intensity $\mu$ is varied within the range $[0,0.5]$, and the decoy intensity $\nu$ within $[0,0.02]$, to optimize the key rate. The optimized values of $\mu$ are displayed in Fig. 5. It is worth noting that some curves in Fig. 4 and Fig. 5 have been smoothed, using function fitting, to reduce numerical fluctuations and to better visualize the trends in bit error rates and intensities $\mu$ under different conditions.
@@
As shown in Fig. 3, our analytical key rates exhibit increasing accuracy with larger values of the phase slice parameter $D$ when compared with numerical results. When $D>7$, the analytical results become nearly indistinguishable from those obtained via numerical optimization, indicating that the analytical approximation is sufficiently tight in this regime. This is because when $D$ increases, we get closer to the CPR case, where there is little difference between numerical and analytical results. Conversely, for smaller values of $D$, the protocol remains functional but achieves secure key exchange only over shorter distances, highlighting the trade-off between implementation simplicity and performance, due to looser bounds on some parameters, such as $e^{b,\mu}_{y,1}$ shown in Fig. 4.
@@
The bit error rate results presented in Fig. 4 help explain the observed differences in key rate performance. The bit error rate $e^{b,\mu}_{y,1}$ determines the bound on $e^{p,\mu}_{x,1}$ in the key rate in Eq. 11, and itself is bounded via Eq. 13. From Eq. 14, we find that the fidelity $F_{\nu 0}$ decreases as $D$ decreases, leading to a looser bound in Eq. 13. Although both methods use the same constraint, numerical methods optimize over the entire range of variables while analytical methods directly pick the worst case for some individual variables, which leads to looser upper bounds in the analytical methods than in the numerical results. For sufficiently large values of $D$, the analytical bounds closely follow the numerical bounds across the entire communication distance. However, as the number of phase slices decreases, the error rates increase rapidly, and the gap between the analytical and numerical bounds becomes more pronounced. This behavior highlights the impact of phase slice resolution on the accuracy of the analytical estimation.
@@
The performance of the DPR BB84 protocol is also closely tied to the choice of signal intensity $\mu$, which directly influences both the secret key rate and the bit error rate. In Fig. 5, when the number of phase slices is $D=10$, the optimized intensity $\mu$, obtained through numerical methods, starts at approximately 0.47. This value is consistent with the findings in [cao2015discrete], where a typical signal intensity of $\mu=0.5$ was reported. As the number of phase slices decreases or the communication distance increases, the optimized $\mu$ correspondingly decreases in order to maintain a positive secret key rate. This behavior can be explained as follows. When $D$ is small, the phase-randomized states become more distinguishable, which compromises the fidelity between the prepared states. To ensure the fidelity in Eq. 7 remains high, we must reduce $\mu$; otherwise, the estimation of the phase error rate $e^{p,\mu}_{x,1}$ in Eq. 16 becomes unreliable. Similarly, at longer communication distances, the channel loss increases significantly, which leads to greater information leakage or error correction cost, quantified by the term $fQ_{x}^{\mu}h_{2}(E_{x}^{\mu})$ in Eq. 11. To mitigate this, the system must again reduce $\mu$, both to suppress multi-photon emissions and to preserve a positive key rate under adverse channel conditions.
@@
Figure 6: Secret key generation rates for the MDI-QKD protocol using numerical (solid lines) and analytical (dashed) methods. The mean number of photons for the signal and decoy states are optimized to get the maximum key rate. The best performing curve is based on the vacuum+weak decoy states and CPR. The remaining curves are generated when considering $D$ phase slices and vacuum+weak decoy states. The parameters used are give in Table I [gobby2004quantum, cao2020discrete].
@@
Figure 7: Bit error rate, $e_{y,11}^{b,\mu\mu}$, for the DPR MDI QKD protocol using numerical (solid) and analytical (dashed) methods. Results for the CPR case are shown as well. For clarity, the curves corresponding to $D<14$ are smoothed to better display the relationship between the bit error rate and either the phase slice number or the communication distance.
@@
Figure 8: The optimum intensity $\mu$ for different values of $D$ versus distance in the DPR MDI QKD protocol. All curves are smoothed to mask numerical fluctuations.
@@
Figure 6 shows the secret key generation rates for the DPR MDI QKD using the vacuum+weak decoy state technique. The corresponding bit error rates are shown in Fig. 7 and the trend of the optimized $\mu$ is depicted in Fig. 8. The key rates are computed using two approaches: the numerical optimization method presented in Section III-B, Section III-A, Approximate Formulation (solid lines), and our analytically derived solutions described in Section IV-B (dashed lines). The simulation parameters are shown in Table I. To determine the optimal key rate, we optimize the intensity $\mu$ within the range $[0,0.4]$ and the decoy intensity $\nu$ within $[0,0.02]$. It is worth noting that some curves in Fig. 7 and Fig. 8 have been smoothed, using function fitting, to better visualize the trends of bit error rates and intensity $\mu$ under different conditions.
As shown in Fig. 6, once the number of discrete phases $D$ reaches 14, the performance gap between the DPR protocol and the fully randomized case becomes negligible, which is consistent with the result in [cao2020discrete]. Our analytical solutions exhibit strong agreement with the numerically optimized results for both the CPR and DPR cases with phase slice numbers $D>10$. Although numerical optimization provides accurate performance estimates, it is computationally intensive and may hinder real-time parameter estimation. In contrast, our analytical bounds offer a fast and practical alternative for determining system parameters and the required level of privacy amplification.
@@
Fig. 7 shows that our analytical bit error rate results closely match the numerical ones. The discrepancy observed at 0 km is attributed to the formulation in Eq. 19, where two $\epsilon$ parameters are introduced during the variable transformation. Although both methods are subject to the same constraints, analytical bounds always select the worst-case value for each variable independently. In contrast, numerical bounds can become tight when the full set of variables is jointly considered. This discrepancy leads to an initial mismatch.
@@
We also compare the optimized $\mu$ with respect to different phase slices and communication distances in Fig. 8. As shown in Fig. 8, the maximum value of the optimized signal intensity $\mu$ is approximately 0.37 in the CPR case. This value gradually decreases as the number of phase slices is reduced or as the communication distance increases, which is caused by the same reason as we discussed in Fig. 5.

### v2

@@ IV-B Analytical bounds for MDI QKD with DPR
In the DPR case, a lower bound on the key rate of the MDI-QKD protocol is given by Eq. 17. The variables we need to bound are $Y_{\gamma,kk}^{\mu\mu}=Q_{\gamma,kk}^{\mu\mu}/(p_{k}^{\mu}p_{k}^{\mu})$ and $e_{x,kk}^{b,\mu\mu}$, for $k\in\{0,1\}$, from which $e_{z,kk}^{p,\mu\mu}$ can be bounded.
@@
$W_{x,00}^{\mu\mu}$ $\leq W_{x,00}^{\mu\mu U}\equiv\max(Q_{x}^{00}E_{x}^{00}+2\epsilon_{\mu 0},0.5),$ (26)
where $Q_{x}^{00}$ and $Q_{x}^{00}E_{x}^{00}$ are observables, and $W_{\gamma,k}^{\alpha\beta}=e_{\gamma,k}^{b,\alpha\beta}Y_{\gamma,k}^{\alpha\beta}$. An upper bound for the bit error rate is $e^{b,\mu\mu}_{x,00}\leq e^{b,\mu\mu U}_{x,00}=W_{x,00}^{\mu\mu U}/Y_{x,00}^{\mu\mu L}$.
In terms of bounds on $Y_{\gamma,11}^{\mu\mu}$ and $W_{x,11}^{\mu\mu}$, we use Eq. 19 and some mathematical inequalities in Appendix B to obtain
@@
can be calculated from observed values; $G=\text{max}(A,B,C)$, where
@@
$\epsilon_{s}=$ $e^{\alpha+\beta}(\sum_{a=2}^{D-1}p_{a}^{\alpha}p_{1}^{\beta}(\epsilon_{\alpha\mu}+\epsilon_{\beta\mu})+\sum_{b=2}^{D-1}p_{1}^{\alpha}p_{b}^{\beta}(\epsilon_{\alpha\mu}+\epsilon_{\beta\mu})$
$+\sum_{a=2}^{D-1}\sum_{b=2}^{D-1}p_{a}^{\alpha}p_{b}^{\beta}(\epsilon_{\alpha\mu}+\epsilon_{\beta\mu})),$
@@
Next, the upper bound on $W_{x,11}^{\mu\mu}$ is calculated from
$\begin{split}W_{x,11}^{\mu\mu}\leq W_{x,11}^{\mu\mu U}\equiv\frac{e^{\alpha+\beta}Q_{x}^{\alpha\beta}E_{x}^{\alpha\beta}-T_{2}+\epsilon_{\bar{s}}}{e^{\alpha+\beta}p_{1}^{\alpha}p_{1}^{\beta}},\end{split}$ (31)
@@
$T_{2}=e^{\alpha+\beta}(p_{0}^{\alpha}Q_{x}^{0\beta}E_{x}^{0\beta}+p_{0}^{\beta}Q_{x}^{\alpha 0}E_{x}^{\alpha 0}-p_{0}^{\alpha}p_{0}^{\beta}Q_{x}^{00}E_{x}^{00}).$ (32)
@@ V Simulation Results and Discussion
In our simulations, we consider the scenario in which no eavesdropper (Eve) is present, but system parameters and imperfections determine the observed values in an experiment. In our simulations of the BB84 protocol, the total gain and the corresponding QBER, for an intensity $\alpha$, are, respectively, assumed to be [21]
$Q_{\gamma}^{\alpha}$ $=1-e^{-\eta\alpha}(1-p_{d})^{2},$
@@
where $p_{d}$ is the dark count rate per pulse, $\eta$ represents the overall channel transmittance, $e_{0}=0.5$ is the error rate associated with vacuum states, and $e_{d}$ denotes the phase misalignment error rate. The nominal values for the simulation parameters are shown in Table I [6, 2], where the optical switch insertion loss is included in the detector efficiency.
In our simulations of the MDI QKD protocol, the total gain and the corresponding QBER when Alice sends states with intensity $\alpha$ and Bob sends states with intensity $\beta$ are, respectively, taken as [17]
$Q_{z}^{\alpha\beta}=$ $Q_{C}+Q_{E},$
$Q_{z}^{\alpha\beta}E_{z}^{\alpha\beta}=$ $e_{d}Q_{C}+(1-e_{d})Q_{E},$
$Q_{x}^{\alpha\beta}=$ $2y^{2}[1+2y^{2}-4yI_{0}(x)+I_{0}(2x)],$
$Q_{x}^{\alpha\beta}E_{x}^{\alpha\beta}=$ $e_{0}Q_{x}^{\alpha\beta}-2(e_{0}-e_{d})y^{2}[I_{0}(2x)-1],$ (34)
where
$Q_{C}=$ $2(1-p_{d})^{2}e^{-\mu^{\prime}/2}(1-(1-p_{d})e^{-\eta\alpha/2})$ (35)
$\times(1-(1-p_{d})e^{-\eta\beta/2}),$
$Q_{E}=$ $2p_{d}(1-p_{d})^{2}e^{-\mu^{\prime}/2}[I_{0}(2x)-(1-p_{d})e^{-\mu^{\prime}/2}],$ (36)
$\mu^{\prime}=\eta(\alpha+\beta)$, $y=(1-p_{d})e^{-\mu^{\prime}/4}$ and $I_{0}(x)$ is the modified Bessel function of the first kind. Using the above values as our observables, we can use the techniques introduced in Section III and Section IV to bound the key rate. Note that, in the numerical results presented, we account for the ten highest order terms in Eq. 5 and Eq. 14, which can be shown to be of sufficiently high precision for the regimes of operation considered.
Figure 3: Secret key generation rates for the DPR BB84 protocol using numerical (solid lines) and analytical (dashed) methods. The mean number of photons for the signal and decoy states are optimized to get the maximum key rate. The best performing curve is based on the vacuum+weak decoy states and CPR. The remaining curves are generated when considering $D$ phase slices and vacuum+weak decoy states.
Figure 4: The upper bound $e_{x,1}^{b,\mu U}$ on the bit error rate for the DPR BB84 protocol using numerical (solid) and analytical (dashed) methods. Results for the CPR case are shown as well.
Figure 5: The optimized value of $\mu$ for different number of phase slices ranging from $D=5$ to $D=10$, versus distance in the DPR BB84 protocol. A CPR related curve is provided for comparison. The numerical results are given in solid lines while analytical results are given in dashed lines.
@@
Figure 3 presents the secret key rates for the vacuum+weak decoy BB84 protocol, using the numerical method (solid lines) described in Section III-A, by solving the optimization problem in Section II-C, Approximate Formulation, and the analytical method (dashed lines) outlined in Section IV-A. The corresponding upper bound on the $X$-basis bit error rate, $e_{x,1}^{b,\mu U}$, is shown in Fig. 4. The simulation parameters are shown in Table I [6, 2]. During the optimization process for both numerical and analytical methods, the intensity $\mu$ is varied within the range $[0,0.5]$, and the decoy intensity $\nu$ within $[0,0.02]$, to optimize the key rate. The optimized values of $\mu$ are displayed in Fig. 5.
@@
As shown in Fig. 3, our analytical key rates exhibit increasing accuracy with larger values of the phase slice parameter $D$ when compared with numerical results. When $D>7$, the analytical results become nearly indistinguishable from those obtained via numerical optimization, indicating that the analytical approximation is sufficiently tight in this regime. This is because when $D$ increases, we get closer to the CPR case, where there is little difference between numerical and analytical results. Conversely, for smaller values of $D$, the protocol remains functional but achieves secure key exchange only over shorter distances, highlighting the trade-off between implementation simplicity and performance, due to looser bounds on some parameters, such as $e^{b,\mu}_{x,1}$ shown in Fig. 4.
@@
The upper bound results on the $X$-basis bit error rate, presented in Fig. 4, can help explain the observed differences in key rate performance. The bit error rate $e^{b,\mu}_{x,1}$ determines the bound on $e^{p,\mu}_{z,1}$ via Eq. 16. The bound on $e^{b,\mu}_{x,1}$ is obtained via Eq. 13 and Eq. 14. From Eq. 14, we find that the fidelity $F_{\nu 0}$ decreases as $D$ decreases, leading to a looser bound in Eq. 13 for $e^{b,\mu}_{x,1}$. Although both methods use the same constraints, numerical methods optimize over the entire range of variables while analytical methods directly pick the worst case for some individual variables, which leads to looser upper bounds in the analytical methods than in the numerical results. For sufficiently large values of $D$, the analytical bounds closely follow the numerical bounds across the entire communication distance. However, as the number of phase slices decreases, the error rates increase rapidly, and the gap between the analytical and numerical bounds becomes more pronounced. This behavior highlights the impact of phase slice resolution on the accuracy of the analytical estimation.
@@
The performance of the DPR BB84 protocol is also closely tied to the choice of signal intensity $\mu$, which directly influences both the secret key rate and the bit error rate. In Fig. 5, when the number of phase slices is $D=10$, the optimized intensity $\mu$, obtained through numerical methods, starts at approximately 0.49. As the number of phase slices decreases or the communication distance increases, the optimized $\mu$ correspondingly decreases in order to maintain a positive secret key rate. This behavior can be explained as follows. When $D$ is small, the phase-randomized states become more distinguishable, which compromises the fidelity between the prepared states. To ensure the fidelity in Eq. 7 remains high, we must reduce $\mu$; otherwise, the estimation of the phase error rate $e^{p,\mu}_{z,1}$ in Eq. 16 becomes unreliable. Similarly, at longer communication distances, the channel loss increases significantly, which leads to greater information leakage or error correction cost, quantified by the term $fQ_{z}^{\mu}h_{2}(E_{z}^{\mu})$ in Eq. 11. To mitigate this, the system must again reduce $\mu$, both to suppress multi-photon emissions and to preserve a positive key rate under adverse channel conditions.
@@
Figure 6: Secret key generation rates for the MDI-QKD protocol using numerical (solid lines) and analytical (dashed) methods. The mean number of photons for the signal and decoy states are optimized to get the maximum key rate. The best performing curve is based on the vacuum+weak decoy states and CPR. The remaining curves are generated when considering $D$ phase slices and vacuum+weak decoy states.
@@
Figure 7: The upper bound on the bit error rate $e_{x,11}^{b,\mu\mu}$, for the DPR MDI QKD protocol using numerical (solid) and analytical (dashed) methods. Results for the CPR case are shown as well.
@@
Figure 8: The optimum intensity $\mu$ for different values of $D$ versus distance in the DPR MDI QKD protocol. The numerical results are shown in solid lines while analytical results are in dashed lines.
@@
Figure 6 shows the secret key generation rates for the DPR MDI QKD using the vacuum+weak decoy state technique. The corresponding upper bound on $e_{x,11}^{b,\mu\mu}$ is shown in Fig. 7 and the optimized $\mu$ is depicted in Fig. 8. The key rates are computed using two approaches: the numerical optimization method presented in Section III-B, Section III-A, Approximate Formulation (solid lines), and our analytically derived solutions described in Section IV-B (dashed lines). The simulation parameters are shown in Table I. To determine the optimal key rate, we optimize the intensity $\mu$ within the range $[0,0.4]$ and the decoy intensity $\nu$ within $[0,0.02]$.
As shown in Fig. 6, once the number of discrete phases $D$ reaches 14, the performance gap between the DPR protocol and the fully randomized case becomes negligible, which is consistent with the result in [3]. Our analytical solutions exhibit strong agreement with the numerically optimized results for both the CPR and DPR cases with phase slice numbers $D>10$. Note that, in Fig. 6, we need more phase slices for MDI QKD than for BB84 to approach the performance offered by CPR. This is due to two main factors. First, in MDI QKD, it is necessary to ensure that the fidelity of the states in Eq. 9 from both parties approaches unity, which demands more phase slices compared to the single-side fidelity in Eq. 7. Secondly, in Eq. 19, obtaining a tight bound requires additional phase slices because the expression involves two $\epsilon$ terms, whereas Eq. 13 contains only one.
@@
Fig. 7 shows that, when $D$ is sufficiently large, our analytical results for the upper bounds on the bit error rate closely match that of the numerical ones. The discrepancy observed at 0 km is attributed to the formulation in Eq. 19, where two $\epsilon$ terms contribute to our analytical bound. Although both methods are subject to the same constraints, analytical bounds always select the worst-case value for each variable independently. In contrast, numerical bounds can become tight when the full set of variables is jointly considered. This discrepancy leads to an initial mismatch.
@@
We also compare the optimized $\mu$ with respect to different phase slices and communication distances in Fig. 8. As shown in Fig. 8, the maximum value of the optimized signal intensity $\mu$ is approximately 0.37 in the CPR case. This value gradually decreases as the number of phase slices is reduced or as the communication distance increases, which is caused by the same reasons as we discussed in Fig. 5.
Overall, in all results presented, we observe a speed up in processing time, when we use our analytical bounds. Although numerical optimization provides accurate performance estimates, it is computationally intensive and may hinder real-time parameter estimation. In contrast, our analytical bounds offer a fast and practical alternative for determining system parameters and the required level of privacy amplification.